a) Assume that Peabody uses a traditional costing system, with overhead allocated based on direct labor hours.  

Compute the total cost per unit of each of the three products.

b) Instead assume that Peabody uses an ABC system with three cost drivers: machine hours, inspections, and setups.
Manufacturing overhead cost of $1,300,000 is due to setups, and $130,000 to inspections, with the rest allocated to machine hours.
Compute the total cost per unit of the three products.
The following table contains per unit data for the three products.
A B C
Raw materials ($) 156 179 185
Direct labor ($) 84 98 70
Machine hours 2.5 2.6 2.1
The following table contains aggregate data for the year for the three products.
A B C
Budgeted production (units) 25,000 18,000 22,000
Number of inspections 250 180 220
Number of setups 100 60 90
Other information:
Wage rate is $28/hour
Budgeted overhead for the year is $13,896,000.
al cost per unit of the three products.
Answer #1
Please give positive ratings so I can keep answering. It would help me a lot. Please comment if you have any query. Thanks!
Peabody
Calculation of labor hours Product A Product B Product C Total Note
Direct Labor cost per unit                                84.00                   98.00              70.00 A
Wage rate per hour                                28.00                   28.00              28.00 B
Labor hour per unit                                  3.00                      3.50                2.50 C=A/B
Budgeted production (units)                        25,000.00            18,000.00      22,000.00 D
Labor hours used                        75,000.00           63,000.00      55,000.00        193,000.00 E=C*D
Answer a
Calculation of Predetermined OH Rate Total Note
Total Direct Labor Hours        193,000.00 See E
Total Manufacturing overhead 13,896,000.00 F
Predetermined OH Rate                  72.00 G=F/E
Allocated OH as per Predetermined OH Rate Product A Product B Product C Total Note
Total Direct Labor Hours                                  3.00                      3.50                2.50 See C
Predetermined OH Rate                                72.00                   72.00              72.00 See G
Allocated OH                             216.00                 252.00            180.00                648.00 H=G*C
Product cost Product A Product B Product C Note
Direct Materials                             156.00                 179.00            185.00
Direct Labor                                84.00                   98.00              70.00
Allocated OH per unit                             216.00                 252.00            180.00 See H
Product cost                             456.00                 529.00            435.00 I
Answer b
Calculation of machine hours Product A Product B Product C Total Note
Machine hour per unit                                  2.50                      2.60                2.10 J
Budgeted production (units)                        25,000.00            18,000.00      22,000.00 See D
Machine hours used                        62,500.00           46,800.00      46,200.00        155,500.00 K=J*D
L M
Calculation of ABC rates for overhead Activity used N=L/M
Activity cost pool Cost Driver Cost Product A Product B Product C Total ABC rates
Machine hours used Machine hours    12,466,000.00      62,500.00           46,800.00      46,200.00       155,500.00                   80.17
Number of inspections Number of inspections         130,000.00            250.00                180.00            220.00               650.00                 200.00
Number of setups Number of setups      1,300,000.00            100.00                  60.00              90.00               250.00              5,200.00
Total Cost assigned 13,896,000.00
See N O P=N*O Q R=N*Q S T=N*S
Allocation table Product A
Product B
Product C
Activity cost pool Activity Measure ABC Rates Activity used Cost assigned Activity used Cost assigned Activity used Cost assigned
Machine hours used Machine hours                   80.17      62,500.00     5,010,450.16      46,800.00    3,751,825.08            46,200.00    3,703,724.76
Number of inspections Number of inspections                 200.00            250.00           50,000.00            180.00         36,000.00                 220.00         44,000.00
Number of setups Number of setups              5,200.00            100.00        520,000.00              60.00       312,000.00                   90.00       468,000.00
Total Cost assigned     5,580,450.16 4,099,825.08 4,215,724.76 U
Budgeted production (units)           25,000.00         18,000.00         22,000.00
Cost assigned per unit                223.22               227.77               191.62
Product cost- ABC System Product A Product B Product C Note
Direct Materials                             156.00                 179.00            185.00
Direct Labor                                84.00                   98.00              70.00
Allocated OH per unit                             223.22                 227.77            191.62 See U
Product cost                             463.22                 504.77            446.62